# HeSum: a Novel Dataset for Abstractive Text Summarization in Hebrew

### v1

@@ 1 Introduction
Recent advances with large language models (LLMs, Brown et al., 2020; Chowdhery et al., 2023) demonstrate impressive capabilies, encompassing diverse tasks such as natural language (NL) understanding and reasoning, including classification tasks such as commonsense reasoning Bisk et al. (2020) and sentiment analysis Liang et al. (2022), as well as generative tasks like summarization and dialogue systems Cohen et al. (2022). However, these impressive achievements are primarily demonstrated for the English language. Our understanding of how these models perform on low-resource languages is limited, as current evaluations are primarily focused on languages with abundant data (Ahuja et al., 2023; Lai et al., 2023).
@@ 3.2 Linguistic Analysis
3.3 Summerization Intrinsic Analysis
@@ 4.2 Results
Table 3 summarizes the quantitative evaluation results. While mLongT5 consistently achieved higher ROUGE scores, a metric focused on surface-level similarity, GPT-based models exhibited superior performance in BertScore, a semantic similarity metric, and in human evaluation scores that assess coherence and completeness. The consistently higher ROUGE scores of mLongT5 might be partially attributed to limitations in the ROUGE metric itself. ROUGE scores favor summaries that closely mimic the source text, even if they lack originality or fluency. Additionally, n-gram based metrics like ROUGE may discount grammatically correct sentences that convey the required meaning even with morphological or lexical word variations, or changes in word order, compared to the source text.
Furthermore, when comparing ROUGE to human evaluation we found negative correlation of human evaluation with ROUGE scores. We computed Pearson correlation coefficients (PCC) and found the coefficients to be around -0.16 with highly statistically significant p-values (less than $2.39\times 10^{-5}$), indicating that higher ROUGE scores do not in actuality correspond to human evaluations of good summaries. Similarly, using Kendall’s $\tau$ correlations resulted in negative values. Further research is needed for developing automatic summarization metrics that correlate with human scores.
@@ 5 Conclusion
Metrics based on n-gram matching, such as BLEU (Papineni et al., 2002), ROUGE Lin (2004), and Meteor Denkowski and Lavie (2014), are commonly used for evaluating summarization quality in English. However, these metrics can be problematic when applied to Hebrew text. Hebrew allows for more flexible word order compared to English. Additionally, its morphological richness entails that the same concept can be expressed in multiple ways due to variations in prefixes, suffixes, and root conjugations. Furthermore, Hebrew has variations in spelling words due to missing vowels (Ktiv haser vs. Ktiv male). These factors can lead to ngram-based metrics overlooking grammatically correct sentences in the generated summary that convey the full meaning even though they show slight differences. Our findings in Section 4.2, which demonstrate a negative correlation between ROUGE scores and human evaluation scores, highlight the limitations of ROUGE evaluation in the context of Hebrew summarization.

### v2

@@ 1 Introduction
Recent advances with large language models (LLMs, Brown et al., 2020; Chowdhery et al., 2023) demonstrate impressive capabilities, encompassing diverse tasks such as natural language (NL) understanding and reasoning, including classification tasks such as commonsense reasoning Bisk et al. (2020) and sentiment analysis Liang et al. (2022), as well as generative tasks like summarization and dialogue systems Cohen et al. (2022). However, these impressive achievements are primarily demonstrated for the English language. Our understanding of how these models perform on low-resource languages is limited, as current evaluations are primarily focused on languages with abundant data (Ahuja et al., 2023; Lai et al., 2023).
@@ 3.2 Linguistic Analysis
3.3 Summarization Intrinsic Analysis
@@ 4.2 Results
Table 3 summarizes the quantitative evaluation results. While mLongT5 consistently achieved higher ROUGE scores, a metric focused on surface-level similarity, GPT-based models exhibited superior performance in BertScore, a semantic similarity metric, and in human evaluation scores that assess coherence and completeness. The consistently higher ROUGE scores of mLongT5 might be partially attributed to limitations in the ROUGE metric itself. ROUGE scores favor summaries that closely mimic the source text, even if they lack originality or fluency. Additionally, n-gram-based metrics like ROUGE may discount grammatically correct sentences that convey the required meaning even with morphological or lexical word variations, or changes in word order, compared to the source text.
Furthermore, when comparing ROUGE to human evaluation we found a negative correlation of human evaluation with ROUGE scores. We computed Pearson correlation coefficients (PCC) and found the coefficients to be around -0.16 with highly statistically significant p-values (less than $2.39\times 10^{-5}$), indicating that higher ROUGE scores do not in actuality correspond to human evaluations of good summaries. Similarly, using Kendall’s $\tau$ correlations resulted in negative values. Further research is needed for developing automatic summarization metrics that correlate with human scores.
@@ 5 Conclusion
Metrics based on n-gram matching, such as BLEU (Papineni et al., 2002), ROUGE Lin (2004), and Meteor Denkowski and Lavie (2014), are commonly used for evaluating summarization quality in English. However, these metrics can be problematic when applied to Hebrew text. Hebrew allows for more flexible word order compared to English. Additionally, its morphological richness entails that the same concept can be expressed in multiple ways due to variations in prefixes, suffixes, and root conjugations. Furthermore, Hebrew has variations in spelling words due to missing vowels (Ktiv haser vs. Ktiv male). These factors can lead to n-gram-based metrics overlooking grammatically correct sentences in the generated summary that convey the full meaning even though they show slight differences. Our findings in Section 4.2, which demonstrate a negative correlation between ROUGE scores and human evaluation scores, highlight the limitations of ROUGE evaluation in the context of Hebrew summarization.